# The DESI DR1/DR2 evidence for dynamical dark energy is biased by low-redshift supernovae

### v4

@@ Abstract
Recently, a $3\sim 4\sigma$ preference for dynamical dark energy has been reported by the Dark Energy Spectroscopic Instrument (DESI) collaboration, which has inspired hot debates on new physics or systematics. In this letter, we reveal that this preference is significantly biased by an external low-redshift supernova (low-$z$ SN) sample, which was combined with the Dark Energy Survey SN program (DES-SN) in their Year-Five data release (DESY5). Using the intercept in the SN magnitude-distance relation as a diagnostic for systematics, we find not only large dispersions but also a large discrepancy in the low-$z$ SN sample when compared to the high-$z$ DES-SN sample within the single DESY5 compilation. Correcting for this low-$z$ systematics with or without including the cosmic microwave background data can largely reduce the preference for dynamical DE to be less than $2\sigma$. Therefore, the DESI preference for dynamical DE is biased by some unknown systematics in the low-$z$ SN sample.
cosmology, supernovae, dark energy
show]schwang@itp.ac.cn
1 Introduction
@@
The standard Lambda-cold-dark-matter ($\Lambda$CDM) model, consisting of radiation, baryonic matter, cold dark matter, and a cosmological constant $\Lambda$ in a spatially flat, isotropic and homogeneous Universe with primordial perturbations provided by a nearly scale-invariant adiabatic Gaussian random scalar field, has passed a wide variety of cosmological tests (e.g., A. G. Riess et al., 1998; D. J. Eisenstein et al., 2005; S. Raghunathan et al., 2019; N. Aghanim et al., 2020; S. Alam et al., 2021; T. M. C. Abbott et al., 2021; M. S. Madhavacheril et al., 2024; C. L. Bennett et al., 1996; G. Hinshaw et al., 2013). Despite its great success over past two decades, a few tensions (see, L. Perivolaropoulos & F. Skara, 2022, for a recent review) including the $\sim 5\sigma$ Hubble tension (J. L. Bernal et al., 2016; L. Verde et al., 2019; L. Knox & M. Millea, 2020; A. G. Riess, 2019; W. L. Freedman, 2021), a $\sim 3\sigma$ $S_{8}$ tension (E. Di Valentino et al., 2021; E. Abdalla et al., 2022), and a $\sim 4\sigma$ $\gamma$ tension (N.-M. Nguyen et al., 2023) as well as a $\sim 3\sigma$ $\delta H_{0}$ tension (W.-W. Yu et al., 2022) seem to breakdown this concordant model. In particular, whether the cosmological constant $\Lambda$ should be replaced by a dynamical dark energy (DE) to drive the late-time acceleration would call for special cautions.
@@
As one of the stage-III surveys, the Dark Energy Survey (DES) explores the properties of dark matter and DE with unprecedented precision and accuracy through multiple probes like Type Ia supernovae (SNe Ia), weak lensings, large-scale structures, galaxy clusters (T. Abbott et al., 2005; J. P. Bernstein et al., 2012; T. Abbott et al., 2016). As the most updated Stage IV survey, the recent release of the Dark Energy Spectroscopic Instrument (DESI) for baryon acoustic oscillations (BAO) (A. G. Adame et al., 2025a, b, c) together with full 5-year photometrically classified 1635 SNe Ia (DES-SN), after combined with previous external 194 low-$z$ SNe Ia (low-$z$ SN) in the DES Year-Five Data Release (DESY5) (T. M. C. Abbott et al., 2024; M. Vincenzi et al., 2024), have triggered heated debates on the existence of dynamical DE. Specifically, the joint constraint of cosmic microwave background (CMB) Planck data, DESI DR1 (Y1 BAO measurements), and DESY5 (low-$z$ SN+DES-SN) compilation surprisingly favors a dynamical DE with phantom crossing, showing a $3.9\sigma$ tension with the $\Lambda$CDM model (A. G. Adame et al., 2025a, 2024). A similar but weaker preference ($\sim 2\sigma$: C.-G. Park et al., 2024a) of dynamic DE is also achieved without the participation of DESI DR1 and DESY5. These hints for new physics have attracted significant interests in the community with subsequent numerous theoretical studies on different kinds of dynamical DE models (e.g., Y. Tada & T. Terada, 2024; K. V. Berghaus et al., 2024; D. Shlivko & P. J. Steinhardt, 2024; O. F. Ramadan et al., 2024; S. Bhattacharya et al., 2024; W. Giarè et al., 2024; I. J. Allali et al., 2024; K. S. Croker et al., 2024; Y. Yang et al., 2024; C. Escamilla-Rivera & R. Sandoval-Orozco, 2024; W. Yin, 2024; G. Gu et al., 2024; X. Wang et al., 2024; M. Cortês & A. R. Liddle, 2024; E. O. Colgáin et al., 2024a; Y. Carloni et al., 2024; K. Lodha et al., 2024; R. Calderon et al., 2024; O. Luongo & M. Muccino, 2024; P. Mukherjee & A. A. Sen, 2024; B. R. Dinda, 2024; R. Camilleri et al., 2024; T.-N. Li et al., 2024; Q. Gao et al., 2024; R. Fikri et al., 2024; T.-N. Li et al., 2025).
@@
In parallel with theoretical explanations, inconsistencies between recent new measurements (DESI DR1, DESY5) and previous high-quality measurements, such as SDSS BAO (S. Alam et al., 2017, 2021) and PantheonPlus SNe (A. G. Riess et al., 2022; D. Brout et al., 2022a; D. Scolnic et al., 2022; D. Brout et al., 2022b; E. R. Peterson et al., 2021; A. Carr et al., 2022; B. Popovic et al., 2023), have also been widely debated (e.g., C.-G. Park et al., 2024b; C.-G. Park & B. Ratra, 2025; G. Liu et al., 2024; E. O. Colgáin et al., 2024b; E. O. Colgáin & M. M. Sheikh-Jabbari, 2024; Z. Wang et al., 2024; Z. Huang et al., 2024; S. Dhawan et al., 2024; S. Roy Choudhury & T. Okumura, 2024; I. D. Gialamas et al., 2024; G. Efstathiou, 2024; A. Notari et al., 2024; Z. Sakr, 2025a, b). For example, the combination of DESI LRG1 and LRG2 data is found to be the key driver for the preference of dynamical DE (G. Liu et al., 2024; Z. Wang et al., 2024; Z. Huang et al., 2024; A. G. Adame et al., 2025a; E. O. Colgáin et al., 2024a). Nevertheless, these two data exhibit mild tension with SDSS BAO (S. Alam et al., 2021) for unknown reasons. What’s more, by comparing the mean magnitude-distance residuals between PantheonPlus and DESY5, G. Efstathiou (2024) found a $\sim 0.04-0.05$ magnitude (mag.) difference in the overlapping regime, indicating unidentified SNe systematics in DESY5 (see, however, M. Vincenzi et al., 2025). Correcting this offset would reduce the preference for dynamical DE, bringing cosmological constraints of the corrected DESY5 back to $\Lambda$CDM. Therefore, it cannot simply rule out the possibility of potential systematics in DESY5.
@@
The low-$z$ SN sample consists of 194 SNe Ia, including 8 SNe Ia from the Carnegie Supernova Project (CSP: K. Krisciunas et al., 2017), 68 SNe Ia from the Center for Astrophysics (CfA: M. Hicken et al., 2009, 2012), and 118 SNe Ia from the PanSTARRS Foundation Supernova Survey (Foundation: R. J. Foley et al., 2018). Including the low-$z$ SN sample in the DESY5 compilation with the least systematics control in cosmological analyses is controversial in practice. As highlighted by T. M. C. Abbott et al. (2024) & M. Vincenzi et al. (2024), the spectroscopically confirmed low-$z$ SN sample (different from the photometrically classified DES-SN in the observational method) is the most critical source of systematics due to a lack of homogeneity and complete calibrations. Specifically, the nuisance parameters $\alpha,\beta$ used to correct for SN color and stretch dependencies show significant discrepancies between low-$z$ SN and DES-SN samples. Given that the properties and selection effects of the low-$z$ sample still lack sufficient understanding, additional systematics checks are necessary.
@@
In this letter, we first cross-compare DESI DR1 and DESY5 SNe with the previous high-quality SDSS eBOSS BAO and PantheonPlus SNe, and confirm that the key driver of dynamical DE is the additional external low-$z$ SN sample in forming the DESY5 compilation. We then propose to diagnose the systematics of DESY5 and quantify its preference for dynamic DE through the dubbed $a_{B}$ consistency developed by L. Huang et al. (2025, 2024). By performing simple corrections to DESY5, we find the fully de-biased DESY5, after combining with DESI DR1/DR2 regardless of CMB, shows no significant deviation from the $\Lambda$CDM model. Therefore, it is too early to claim dynamical DE.
2 Data and method
@@
As indicated by A. G. Adame et al. (2025a, 2024), different late-time observations seem to have different levels of support for the dynamical DE. To diagnose which dataset contributes the most to the preference for dynamical DE, we take all the high-quality DESI DR1 (A. G. Adame et al., 2025a), DESY5 (low-$z$ SN + DES-SN: T. M. C. Abbott et al., 2024; M. Vincenzi et al., 2024), SDSS eBOSS BAO (S. Alam et al., 2021), PantheonPlus (A. G. Riess et al., 2022; D. Brout et al., 2022a; D. Scolnic et al., 2022; D. Brout et al., 2022b; E. R. Peterson et al., 2021; A. Carr et al., 2022; B. Popovic et al., 2023), and Planck CMB (N. Aghanim et al., 2020) into considerations for completed comparisons. The UNION3 compilation (D. Rubin et al., 2023) is not included in this study as they have not released a Hubble diagram yet.
Similar to PantheonPlus SNe analysis, we construct the distance moduli residual vector of DESY5 SNe as $\Delta\vec{D}=m_{B}^{\mathrm{std}}-M_{B}-\mu_{\mathrm{model}}$, where standardized apparent magnitudes $m_{B}^{\mathrm{std}}$ have already corrected for stretch, color, simulation bias (selection bias, Malmquist bias, and light curve fitting bias), and mass step effects. The likelihood is defined as $-2\ln(\mathcal{L})=\Delta\vec{D^{T}}C^{-1}_{\mathrm{stat+syst}}\Delta\vec{D}$. We adopt the commonly used built-in Planck CMB (plik TTTEEE + lowl + lowE + CMB lensing), SDSS eBOSS BAO (eight BAO-only data points), and PantheonPlus (local SNe at $z<0.01$ are discarded) in MontePython (B. Audren et al., 2013), along with extra self-written DESI BAO and DESY5 likelihoods. In what follows, we perform the Monte Carlo Markov Chain (MCMC) on the $w_{0}w_{a}$ model with Planck CMB + arbitrarily different combinations of BAO and SNe to locate the possible systematics direction.
3 Analysis
3.1 Systematics identification
Figure 2: The $a_{B}$ diagnosis: the distributions of $-5a_{B}$ for low-$z$, DES-SN, and PantheonPlus SNe (only 1365 SNe with $z\geq 0.0233$ are considered). The fiducial cosmology is assumed from the best-fit Planck $\Lambda$CDM with $\Omega_{m}=0.315$. Note that the discrepancy in $-5a_{B}$ between low-$z$ and DES-SN samples is independent of fiducial $\Omega_{m}$ values. The top panel is the redshift distribution for the number of SNe while the right panel is the probability density distributions of $-5a_{B}$ for different SN samples.
@@ 3.1 Systematics identification
To further separate the interference from different BAO data and reflect the underlying preference for dynamical DE from different SNe samples, we can solely use Planck CMB to calibrate low-$z$, DES-SN, and their combined (DESY5), respectively. The marginalized $w_{0}-w_{a}$ and $H_{0}-M_{B}$ distributions are shown in the right panel of Fig. 1. The purple $w_{0}-w_{a}$ contour of low-$z$ SN samples shows a significant deviation and thus dominates the preference for dynamical DE in the blue DESY5 contour since the red DES-SN contour alone coincides perfectly with $\Lambda$CDM ($\sim 0.6\sigma$). However, this low-$z$-dominated deviation is likely to be a mirage since its $H_{0}-M_{B}$ contour in the inset shows a rather weak constraint and irregular marginalized distribution, suggesting the multi-sample mixed low-$z$ SNe does not have a consistent absolute magnitude $M_{B}$. Hence, there are substantial systematics in the low-$z$ SNe sample.
3.2 Systematics diagnosis
@@
for both low-$z$ and DES-SN samples, where the dimensionless luminosity distance $\hat{d}_{L}(z_{i})=D_{L}(z)/(c/H_{0})=(1+z_{\mathrm{hel}})\int_{0}^{z_{\mathrm{HD}}}\mathrm{d}z^{\prime}/E(z^{\prime})$ is computed from the best-fit Planck $\Lambda$CDM fiducial model for the reduced Hubble parameter $E(z)^{2}\equiv H(z)^{2}/H_{0}^{2}=\Omega_{m}(1+z)^{3}+(1-\Omega_{m})$. Recall that all light-curve model corrections are used to standardize the apparent magnitude $m_{B,i}^{\mathrm{std}}$, and hence the intercept $-5a_{B}$ as a degeneracy parameter between $H_{0}$ and $M_{B}$ is expected to be a stable value for a single SNe compilation. The gist of $a_{B}$ diagnosis is to check whether the intercept $-5a_{B}$ (a systematics-sensitive and model-sensitive quantity) is consistent among different SNe samples. As any observational systematics in measuring $m_{B,i}^{\mathrm{std}}$ or theoretical systematics in modeling $5\lg\hat{d}_{L}(z_{i})$ induced by either uncounted peculiar velocity effect on the redshift $z_{i}$ measurement or even new physics model in $\hat{d}_{L}(z)$ would cause a deviation in the intercept $-5a_{B,i}=m_{B,i}^{\mathrm{std}}-5\lg\hat{d}_{L}(z_{i})$, we can quickly locate the redshift ranges of either systematics or new physics from redshift distributions of $-5a_{B}$.
@@
Detailed $a_{B}$ diagnosis is visualized in Fig. 2. The high-quality DES-SN from a single telescope has the most stable $a_{B}$ distributions among its spanned redshift bins, perfectly agreeing with the Planck $\Lambda$CDM cosmology assumed in Fig. 2. By contrast, the multi-sample mixed low-$z$ SNe shows irregular fluctuations in $-5a_{B}$, and its weighted average is $\sim 0.043$ mag. higher than that of DES-SN, which is qualitatively similar to the previous study by G. Efstathiou (2024) but with the $\sim 0.04$ mag. offset estimated on $\Delta m_{\mathrm{offset}}:=(m_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-m_{\mathrm{DESY5}}^{\mathrm{low}-z})-(m_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-m_{\mathrm{DESY5}}^{\mathrm{high}-z})\approx(-0.05)-(-0.01)=-0.04$ mag. between PantheonPlus and DESY5 compilations in their low- and high-redshift ranges after assuming a constant $M_{B}$. As a response, M. Vincenzi et al. (2025) has explained partially this offset with improvements to the intrinsic scatter model and mass step estimate by reducing the offset from $\sim 0.04$ mag. to $\Delta\mu_{\mathrm{offset}}:=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{low}-z})-(\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z})-(\mu_{\mathrm{DESY5}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=0.019-0.043=-0.024$ mag., which can be further reduced to $-0.008$ mag. after considering a $\sim 0.016$ mag. difference in selection functions as elaborated in Table 1 of M. Vincenzi et al. (2025). However, such a negligible $-0.008$ mag. offset in $\Delta\mu_{\mathrm{offset}}$ is actually achieved between two different SN compilations, leaving an equivalent $0.019+0.008\sim 0.027$ mag. difference between low-$z$ and high-$z$ SNe within the single DESY5 compilation, which is still apparent in the difference of $-5a_{B}$ between low-$z$ and high-$z$ parts of the single DESY5 compilation. Therefore, the key difference of our current study from the previous two analyses (G. Efstathiou, 2024; M. Vincenzi et al., 2025) is that they considered the low/high-redshift difference among two different SN compilations (PantheonPlus and DESY5) while we focus on the low/high-redshift difference within the same SN compilation.

### v5

@@ Abstract
Recently, a $3\sim 4\sigma$ preference for dynamical dark energy has been reported by the Dark Energy Spectroscopic Instrument (DESI) collaboration, which has inspired hot debates on new physics or systematics. In this paper, we reveal that this preference is significantly biased by an external low-redshift supernova (low-$z$ SN) sample, which was combined with the Dark Energy Survey SN program (DES-SN) in their Year-Five data release (DESY5). Using the intercept in the SN magnitude-distance relation as a diagnostic for systematics, we find not only large dispersions but also a large discrepancy in the low-$z$ SN sample when compared to the high-$z$ DES-SN sample within the single DESY5 compilation, in contrast to the uniform behavior found in the PantheonPlus data. Correcting for this low-$z$ systematics with or without including the cosmic microwave background data can largely reduce the preference for dynamical DE to be less than $2\sigma$. Therefore, the DESI preference for dynamical DE is biased by some unknown systematics in the low-$z$ SN sample.
I Introduction
@@
The standard Lambda-cold-dark-matter ($\Lambda$CDM) model, consisting of radiation, baryonic matter, cold dark matter, and a cosmological constant $\Lambda$ in a spatially flat, isotropic and homogeneous Universe with primordial perturbations provided by a nearly scale-invariant adiabatic Gaussian random scalar field, has passed a wide variety of cosmological tests Riess et al. (1998); Eisenstein et al. (2005); Raghunathan et al. (2019); Aghanim et al. (2020); Alam et al. (2021); Abbott et al. (2022); Madhavacheril et al. (2024); Bennett et al. (1996); Hinshaw et al. (2013). Despite its great success over past two decades, a few tensions Perivolaropoulos and Skara (2022) including the $\sim 5\sigma$ Hubble tension Bernal et al. (2016); Verde et al. (2019); Knox and Millea (2020); Riess (2019); Freedman (2021), a $\sim 3\sigma$ $S_{8}$ tension Di Valentino et al. (2021); Abdalla et al. (2022), and a $\sim 4\sigma$ $\gamma$ tension Nguyen et al. (2023) as well as a $\sim 3\sigma$ $\delta H_{0}$ tension Yu et al. (2022) seem to breakdown this concordant model. In particular, whether the cosmological constant $\Lambda$ should be replaced by a dynamical dark energy (DE) to drive the late-time acceleration would call for special cautions.
@@
As one of the stage-III surveys, the Dark Energy Survey (DES) explores the properties of dark matter and DE with unprecedented precision and accuracy through multiple probes like Type Ia supernovae (SNe Ia), weak lensings, large-scale structures, galaxy clusters Abbott et al. (2005); Bernstein et al. (2012); Abbott et al. (2016). As the most updated Stage IV survey, the recent release of the Dark Energy Spectroscopic Instrument (DESI) for baryon acoustic oscillations (BAO) Adame et al. (2025a, b, c) together with full 5-year photometrically classified 1635 SNe Ia (DES-SN), after combined with previous external 194 low-$z$ SNe Ia (low-$z$ SN) in the DES Year-Five Data Release (DESY5) Abbott et al. (2024); Vincenzi et al. (2024), have triggered heated debates on the existence of dynamical DE. Specifically, the joint constraint of cosmic microwave background (CMB) Planck data, DESI DR1 (Y1 BAO measurements), and DESY5 (low-$z$ SN+DES-SN) compilation surprisingly favors a dynamical DE with phantom crossing, showing a $3.9\sigma$ tension with the $\Lambda$CDM model Adame et al. (2025a, d). A similar but weaker preference ($\sim 2\sigma$Park et al. (2024a) of dynamic DE is also achieved without the participation of DESI DR1 and DESY5. These hints for new physics have attracted significant interests in the community with subsequent numerous theoretical studies on different kinds of dynamical DE models, e.g. Refs. Tada and Terada (2024); Berghaus et al. (2024); Shlivko and Steinhardt (2024); Ramadan et al. (2024); Bhattacharya et al. (2024); Giarè et al. (2024); Allali et al. (2025); Croker et al. (2024); Yang et al. (2024); Escamilla-Rivera and Sandoval-Orozco (2024); Yin (2024); Gu et al. (2024); Wang et al. (2024a); Cortês and Liddle (2024); Colgáin et al. (2024); Carloni et al. (2025); Lodha et al. (2025a); Calderon et al. (2024); Luongo and Muccino (2024); Mukherjee and Sen (2024); Dinda (2024); Camilleri et al. (2024); Li et al. (2025a); Gao et al. (2025); Fikri et al. (2024); Li et al. (2025b).
@@
In parallel with theoretical explanations, inconsistencies between recent new measurements (DESI DR1, DESY5) and previous high-quality measurements, such as SDSS BAO Alam et al. (2017, 2021) and PantheonPlus SNe Riess et al. (2022); Brout et al. (2022a); Scolnic et al. (2022); Brout et al. (2022b); Peterson et al. (2022); Carr et al. (2022); Popovic et al. (2023), have also been widely debated Park et al. (2024b); Park and Ratra (2025); Liu et al. (2024); Colgáin et al. (2025); Colgáin and Sheikh-Jabbari (2024); Wang et al. (2024b); Huang et al. (2024); Dhawan et al. (2025); Roy Choudhury and Okumura (2024); Gialamas et al. (2025); Efstathiou (2025); Notari et al. (2025); Sakr (2025). For example, the combination of DESI LRG1 and LRG2 data is found to be the key driver for the preference of dynamical DE Liu et al. (2024); Wang et al. (2024b); Huang et al. (2024); Adame et al. (2025a); Colgáin et al. (2024). Nevertheless, these two data exhibit mild tension with SDSS BAO Alam et al. (2021) for unknown reasons. What’s more, by comparing the mean magnitude-distance residuals between PantheonPlus and DESY5, Ref. Efstathiou (2025) found a $\sim 0.04-0.05$ magnitude (mag.) difference in the overlapping regime, indicating unidentified SNe systematics in DESY5 (see, however, Ref. Vincenzi et al. (2025)). Correcting this offset would reduce the preference for dynamical DE, bringing cosmological constraints of the corrected DESY5 back to $\Lambda$CDM. Therefore, it cannot simply rule out the possibility of potential systematics in DESY5.
@@
The low-$z$ SN sample consists of 194 SNe Ia, including 8 SNe Ia from the Carnegie Supernova Project (CSP) Krisciunas et al. (2017), 68 SNe Ia from the Center for Astrophysics (CfA) Hicken et al. (2009, 2012), and 118 SNe Ia from the PanSTARRS Foundation Supernova Survey (Foundation) Foley et al. (2018). On the other hand, the PantheonPlus sample consists of 741 SNe Ia at low redshift $z<0.1$, 184 of which are in common with the low-$z$ SN sample of DESY5 compilation, that is, 118 SNe from Foundation, 59 SNe from CfA, and 7 SNe from CSP samples. Including the low-$z$ SN sample in the DESY5 compilation with the least systematics control in cosmological analyses is controversial in practice. As highlighted in Refs. Abbott et al. (2024); Vincenzi et al. (2024), the spectroscopically confirmed low-$z$ SN sample (different from the photometrically classified DES-SN in the observational method) is the most critical source of systematics due to a lack of homogeneity and complete calibrations. Specifically, the nuisance parameters $\alpha,\beta$ used to correct for SN color and stretch dependencies show significant discrepancies between low-$z$ SN and DES-SN samples. Given that the properties and selection effects of the low-$z$ sample still lack sufficient understanding, additional systematics checks are necessary.
@@
In this paper, we first cross-compare DESI DR1 and DESY5 SNe with the previous high-quality SDSS eBOSS BAO and PantheonPlus SNe, and confirm that the key driver of dynamical DE is the additional external low-$z$ SN sample in forming the DESY5 compilation. We then propose to diagnose the systematics of DESY5 and quantify its preference for dynamic DE through the dubbed $a_{B}$ consistency developed in Refs. Huang et al. (2025a, b). By performing simple corrections to DESY5, we find the fully de-biased DESY5, after combining with DESI DR1/DR2 regardless of CMB, shows no significant deviation from the $\Lambda$CDM model. Therefore, it is too early to claim dynamical DE.
II Data and method
@@
As indicated in Refs. Adame et al. (2025a, d), different late-time observations seem to have different levels of support for the dynamical DE. To diagnose which dataset contributes the most to the preference for dynamical DE, we take all the high-quality DESI DR1 Adame et al. (2025a), DESY5 (low-$z$ SN + DES-SN) Abbott et al. (2024); Vincenzi et al. (2024), SDSS eBOSS BAO Alam et al. (2021), PantheonPlus Riess et al. (2022); Brout et al. (2022a); Scolnic et al. (2022); Brout et al. (2022b); Peterson et al. (2022); Carr et al. (2022); Popovic et al. (2023), and Planck CMB Aghanim et al. (2020) into considerations for completed comparisons. The UNION3 compilation Rubin et al. (2023) is not included in this study as they have not released a Hubble diagram yet.
Similar to PantheonPlus SNe analysis, we construct the distance moduli residual vector of DESY5 SNe as $\Delta\vec{D}=m_{B}^{\mathrm{std}}-M_{B}-\mu_{\mathrm{model}}$, where standardized apparent magnitudes $m_{B}^{\mathrm{std}}$ have already corrected for stretch, color, simulation bias (selection bias, Malmquist bias, and light curve fitting bias), and mass step effects. The likelihood is defined as $-2\ln(\mathcal{L})=\Delta\vec{D^{T}}C^{-1}_{\mathrm{stat+syst}}\Delta\vec{D}$. We adopt the commonly used built-in Planck CMB (plik TTTEEE + lowl + lowE + CMB lensing), SDSS eBOSS BAO (eight BAO-only data points), and PantheonPlus (local SNe at $z<0.01$ are discarded) in MontePython Audren et al. (2013), along with extra self-written DESI BAO and DESY5 likelihoods. In what follows, we perform the Monte Carlo Markov Chain (MCMC) on the $w_{0}w_{a}$ model with Planck CMB + arbitrarily different combinations of BAO and SNe to locate the possible systematics direction.
III Analysis
III.1 Systematics identification
Figure 2: Top: The $a_{B}$ diagnosis: the distributions of $-5a_{B}$ for low-$z$, DES-SN, and PantheonPlus SNe (only 1365 SNe with $z\geq 0.0233$ are considered). The fiducial cosmology is assumed from the best-fit Planck $\Lambda$CDM with $\Omega_{m}=0.315$. Note that the discrepancy in $-5a_{B}$ between low-$z$ and DES-SN samples is independent of fiducial $\Omega_{m}$ values. The top panel is the redshift distribution for the number of SNe while the right panel is the probability density distributions of $-5a_{B}$ for different SN samples. Bottom: The fiducial cosmology used to calculate $-5a_{B}$ is changed into the best-fit $w_{0}w_{a}$ model from DESI DR2 Abdul Karim et al. (2025a) with $\Omega_{m}=0.352_{-0.018}^{+0.041}$, $w_{0}=-0.48_{-0.17}^{+0.35}$, and $w_{a}<-1.34$.
@@ III.1 Systematics identification
To further separate the interference from different BAO data and reflect the underlying preference for dynamical DE from different SNe samples, we can solely use Planck CMB to calibrate low-$z$, DES-SN, and their combined (DESY5), respectively. The marginalized $w_{0}-w_{a}$ and $H_{0}-M_{B}$ distributions are shown in the right panel of Fig. 1. The purple $w_{0}-w_{a}$ contour of low-$z$ SN samples shows a significant deviation and thus dominates the preference for dynamical DE in the blue DESY5 contour since the red DES-SN contour alone coincides perfectly with $\Lambda$CDM ($\sim 0.6\sigma$). However, this low-$z$-dominated deviation is likely to be a mirage since its $H_{0}-M_{B}$ contour in the inset shows a rather weak constraint and irregular marginalized distribution, suggesting the multi-sample mixed low-$z$ SNe does not have a consistent absolute magnitude $M_{B}$. Hence, there are substantial systematics in the low-$z$ SNe sample, which can be corrected with a method (3) detailed below in the next two subsections, as shown with black dashed contours in both Fig. 1 and its inset.
III.2 Systematics diagnosis
@@
for both low-$z$ and DES-SN samples, where the dimensionless luminosity distance $\hat{d}_{L}(z_{i})=D_{L}(z)/(c/H_{0})=(1+z_{\mathrm{hel}})\int_{0}^{z_{\mathrm{HD}}}\mathrm{d}z^{\prime}/E(z^{\prime})$ is computed from the best-fit Planck $\Lambda$CDM fiducial model for the reduced Hubble parameter $E(z)^{2}\equiv H(z)^{2}/H_{0}^{2}=\Omega_{m}(1+z)^{3}+(1-\Omega_{m})$. Recall that all light-curve model corrections are used to standardize the apparent magnitude $m_{B,i}^{\mathrm{std}}$, and hence the intercept $-5a_{B}$ as a degeneracy parameter between $H_{0}$ and $M_{B}$ is expected to be a stable value for a single SNe compilation. The gist of $a_{B}$ diagnosis is to check whether the intercept $-5a_{B}$ (a systematics-sensitive and model-sensitive quantity) is consistent among different SNe samples. As any observational systematics in measuring $m_{B,i}^{\mathrm{std}}$ or theoretical systematics in modeling $5\lg\hat{d}_{L}(z_{i})$ induced by either uncounted peculiar velocity effect on the redshift $z_{i}$ measurement or even new physics modeling in $\hat{d}_{L}(z)$ would cause a deviation in the intercept $-5a_{B,i}=m_{B,i}^{\mathrm{std}}-5\lg\hat{d}_{L}(z_{i})$, we can quickly locate the redshift ranges of either systematics or new physics from redshift distributions of $-5a_{B}$.
@@
Detailed $a_{B}$ diagnosis is visualized in the top panel of Fig. 2 with a fiducial Planck-$\Lambda$CDM cosmology. The high-quality DES-SN from a single telescope has the most stable $a_{B}$ distributions among its spanned redshift bins, perfectly agreeing with the Planck $\Lambda$CDM cosmology assumed in the top panel of Fig. 2. By contrast, the multi-sample mixed low-$z$ SNe shows irregular fluctuations in $-5a_{B}$, and its weighted average is $\sim 0.043$ mag. higher than that of DES-SN, which is qualitatively similar to the previous study by Efstathiou (2025) but with the $\sim 0.04$ mag. offset estimated on $\Delta m_{\mathrm{offset}}:=(m_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-m_{\mathrm{DESY5}}^{\mathrm{low}-z})-(m_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-m_{\mathrm{DESY5}}^{\mathrm{high}-z})\approx(-0.05)-(-0.01)=-0.04$ mag. between PantheonPlus and DESY5 compilations in their low- and high-redshift ranges after assuming a constant $M_{B}$. As a response, Ref. Vincenzi et al. (2025) has explained partially this offset with improvements to the intrinsic scatter model and mass step estimate by reducing the offset from $\sim 0.04$ mag. to $\Delta\mu_{\mathrm{offset}}:=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{low}-z})-(\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z})-(\mu_{\mathrm{DESY5}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=0.019-0.043=-0.024$ mag., which can be further reduced to $-0.008$ mag. after considering a $\sim 0.016$ mag. difference in selection functions as elaborated in Table 1 of Ref. Vincenzi et al. (2025). However, such a negligible $-0.008$ mag. offset in $\Delta\mu_{\mathrm{offset}}$ is actually achieved between two different SN compilations, leaving an equivalent $0.019+0.008\sim 0.027$ mag. difference between low-$z$ and high-$z$ SNe within the single DESY5 compilation, which is still apparent in the difference of $-5a_{B}$ between low-$z$ and high-$z$ parts of the single DESY5 compilation. Therefore, the key difference of our current study from the previous two analyses Efstathiou (2025); Vincenzi et al. (2025) is that they considered the low/high-redshift difference among two different SN compilations (PantheonPlus and DESY5) while we focus on the low/high-redshift difference within the same SN compilation.